OpenStudy (anonymous):

What are the roots of? PLEASE HELP >>>y=4x^2-20x-24

OpenStudy (blurbendy):

you can start by factoring the four out of each term

OpenStudy (maddyx):

First, factorize the right hand side equation in x by middle term: 4x^2 - 20x -24 = 4x^2 - 24x + 4x - 24 = 4(x+1)(x-6) Now, find the value of x if y was 0 (from the equation: y=4(x+1)(x-6) Those will be your roots